OpenStudy (anonymous):

In Mr . Montgomerys class , there are 8 boys and 12 girls . If Mr . Montgomery selects 1 student from his class at random , what is the probability that the student will be a girl . ?

OpenStudy (hoblos):

the formula says number of favorable casses P= ------------------------- number of all casses

OpenStudy (hoblos):

and here the number of favorable cases is number of girls in class and the number of all cases is the total number of students in class

OpenStudy (hoblos):

can you get it now?

OpenStudy (anonymous):

@hoblos I didn't know this one either... hope this helps me too ( At school never learned about probability !)

OpenStudy (anonymous):

So , will it turn out to be 3/5 ?

OpenStudy (hoblos):

yes you are right @rhonda75 :)

OpenStudy (anonymous):

Is it because , theres 12 girls out of 20 students andthats simplified to 6/10 which simplifies to 3/5 ?

OpenStudy (hoblos):

yes P=12/20 = 3/5

OpenStudy (anonymous):

WooHoo ! I have more questions to ask , can I get you to help me ?

OpenStudy (hoblos):

sure..

OpenStudy (anonymous):

Sean has a spinner and a bag of marbles. The spinner is evenly divided into 8 sections. The bag is filled with 6 different colored marbles. Sean will spin the spinner and randomly select a marble. How many possilbe outcomes does Sean have ?

OpenStudy (anonymous):

Here total outcomes are =8+12=20 favourable (girls)= 12 so P = fav/total= 12/20